# A multi-process imaging of nuclear modifications on parton distributions in proton-nucleus collisions

### v1

@@ Abstract
Nuclear modifications to collinear parton distribution functions are conventionally quantified by the ratios $r^{\textrm{A}}_{i}(x,Q^{2})=f^{\textrm{A,proton}}_{i}(x,Q^{2})/f^{\textrm{proton}}_{i}(x,Q^{2})$. For a given nucleus $A$, these ratios generally depend on the parton momentum fraction $x$, the probing scale $Q^{2}$, and the parton species $i$. Determining these dependencies relies on a global analysis of diverse experimental data. However, in realistic observables, these dependencies are intricately intertwined, making their extraction challenging. In this paper, we propose a novel approach to effectively image the nuclear modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$ at the observable level in proton-nucleus collisions at the Large Hadron Collider. Specifically, through a combined study of $Z$-boson production, $Z$+jet production, and $Z+c$-jet production, we separately enhance signals arising from light-quark, gluon, and heavy-flavor (charm) distributions in nuclei. This enables us to effectively image the $r^{\textrm{A}}_{i}(x,Q^{2})$ for specific parton species. The feasibility of this method is validated through perturbative calculations at next-to-leading order in the strong coupling constant, employing three sets of nuclear PDF parametrizations: EPPS21, nCTEQ15, and TUJU19. Future measurements of these observables are expected to provide more efficient constraints on nuclear PDFs and yield new insights into the detailed partonic structures of nuclei.
@@ I Introduction
Figure 1: Typical measurements of differential cross sections for $Z$-boson production (left panel), $Z\!+\!$jet production (middle panel), and $Z\!+\!c\!-\!$jet production (right panel) in $pp$ collisions at $\sqrt{s}=8$ TeV at the LHC. Measurements are binned according to the rapidity $y^{Z}$ and transverse momentum $p_{T}^{Z}$ for $Z$-boson production CMS:2015hyl , the rapidity $y^{\textrm{jet}}$ and transverse momentum $p_{T}^{\textrm{jet}}$ for $Z\!+\!$jet production ATLAS:2019bsa , and the transverse momentum $p_{T}^{\textrm{jet}}$ for $Z\!+\!c\!-\!$jet production CMS:2017snu . Experimental data are represented by black discs, while NLO predictions computed with MCFM Campbell:1999ah ; Campbell:2011bn ; Campbell:2015qma ; Campbell:2021vlt are depicted by red lines.
The vast amount of data from the Large Hadron Collider (LHC) has revolutionized global analyses over the past decade Klasen:2023uqj ; Klasen:2024xqn , leading to more stringent constraints on nuclear PDFs, notably for gluon distributions Kusina:2017gkz ; Eskola:2019dui ; Eskola:2019bgf ; Kusina:2020dki ; Paakkinen:2022qxn , and advancing neural-network-based analyses that reduce reliance on specific parametrizations AbdulKhalek:2022fyi . However, traditional measurements in proton-proton ($pp$) and proton-nucleus ($p$A) collisions exhibit a less direct mapping to PDFs, and the imaging of $r^{\textrm{A}}_{i}(x,Q^{2})$ achieved in DIS has not been replicated for most LHC processes ALICE:2013snh ; LHCb:2017ygo ; CMS:2018jpl ; ALICE:2023ama ; ATLAS:2024mvt . This makes systematic comparisons of nuclear modifications across different processes challenging. Furthermore, significant discrepancies among widely used nuclear PDF parametrizations in certain regions remain unresolved Klasen:2023uqj ; Klasen:2024xqn .
In a previous study Shen:2021eir , a method to reorganize the differential cross section for dijet production, initially proposed by Ellis and Soper Ellis:1994dg , was extended to $p$A collisions. This approach enables a well-controlled kinematic scan over $x$ and $Q^{2}$ for nuclear PDFs, providing an overall image of the $r^{\textrm{A}}_{i}(x,Q^{2})$ factors Shen:2021eir . However, this image combines contributions from different parton species simultaneously. Given that the relative contributions of various parton species vary across processes, extending this method to other $p$A collision processes is of significant interest. Specifically, once kinematic scans are established for these processes, it becomes possible to enhance signals for specific parton flavors by combining results from different processes.
In this paper, we extend the method from Ref. Shen:2021eir to $Z$-boson, $Z+$jet, and $Z+$charm-jet production in proton-lead ($p$Pb) collisions at the LHC, enabling kinematic scans of nuclear PDFs for each process. The partonic subprocesses at leading order (LO) are listed in Tab. 1.
@@
Table 1: Partonic subprocesses at LO for $Z$-boson production, $Z+$jet production, and $Z+$charm-jet production, respectively.
@@
The paper is organized as follows: In Sec. II, we extend the method from Ref. Shen:2021eir to reorganize the differential cross sections for $Z$-boson, $Z+$jet, and $Z+$charm-jet production in $p$Pb collisions. We then present an imaging of $r^{\textrm{A}}_{i}(x,Q^{2})$ for light quarks using $Z$-boson production. In Sec. III, we design two combinations of multi-process cross sections and show that their nuclear modifications can image $r^{\textrm{A}}_{i}(x,Q^{2})$ for gluon and heavy-flavor (charm) distributions, respectively. A summary and discussion are provided in Sec. IV.
@@ II Reorganizing differential cross section in proton-nucleus collisions
Traditional measurements in $pp$ and $p$A collisions are typically conducted in the form of differential cross sections, often double-differential, binned with variables such as the transverse momentum, rapidity, and invariant mass of the final-state particles Albacete:2013ei ; Albacete:2017qng . For instance, illustrated in Fig. 1 are three representative measurements for $Z$-boson production CMS:2015hyl , $Z+$jet production ATLAS:2019bsa , and $Z+$charm-jet productionCMS:2017snu in $pp$ collisions at the LHC, respectively. In general, such observables may not be ideal for probing the PDFs defined in terms of the parton momentum fraction $x$ and the probing scale $Q^{2}$. For example, in the case of $Z+$jet production, measuring the cross section binned with the jet transverse momentum $p_{T}^{\textrm{jet}}$ and the jet rapidity $y^{\textrm{jet}}$ (as shown in the middle panel of Fig. 1) only constrains the kinematics of the jet, while leaving the kinematics of the $Z$ boson unconstrained. Consequently, even at LO, neither the momentum fraction carried by the initial-state parton nor the physical scale of the process can be precisely deduced from these two measured variables alone.
A well-controlled kinematic scan of the PDFs can be achieved by reorganizing the measured variables to establish a correspondence with $x$ and $Q^{2}$ at LO Shen:2021eir . To constrain the momentum fraction carried by the initial-state nuclear parton in $p$A collisions (with the nucleus moving backward), a variable $X_{B}$ can be introduced Shen:2021eir ; Ellis:1994dg , generally expressed as
@@
Here, the summation of the backward components of the light-cone momenta, $p_{i}^{-}=E_{Ti}e^{-y_{i}}$, is performed over all inclusively produced particles. Specifically, for $Z$-boson production with leptonic decay ($Z\rightarrow l^{+}l^{-}$), the summation includes the di-lepton, while for $Z+$jet or $Z\!+\!c$-jet production, it encompasses both the particles in the jet and the di-lepton from $Z$ decay. With this definition, $X_{B}$ precisely equals the momentum fraction $x_{\textrm{Pb}}$ of the parton from the lead nucleus in $p$Pb collisions at LO accuracy. However, this equivalence may not hold at NLO and beyond.
@@
For $Z\!+\!(c)$jet production, which involves more complex dynamic channels, one can employ both the invariant mass of the $Z$-jet system, $M_{V\!J}$, and the averaged transverse momentum, $p_{T,avg}$, to define a triple differential cross section:
@@
As an initial test, we investigate the $R_{pA}$ factor for the reorganized differential cross section of $Z$-boson production in $p$Pb collisions, as defined in Eq. (2). Since the partonic processes at LO are initiated by a $q\bar{q}$ pair, the $R_{pA}$ factors are expected to primarily reflect the nuclear modifications on the light-quark distributions. In the three panels of Fig. 3, the NLO results for $R_{pA}$ are plotted as a function of $X_{B}$ (blue squares) using three sets of nuclear PDFs. The variable $M_{V}$ is constrained around the $Z$-mass peak, specifically within the range $11666<M_{V}<$ 116 GeV. To contextualize these modifications, the $r^{\textrm{A}}_{i}(x,Q^{2})$ curves for light-quark (solid) and gluon (dotted-dashed) distributions from the corresponding nuclear PDFs are also plotted in each panel of Fig. 3 as a function of $x_{\textrm{Pb}}$, with the probing scale $Q$ set at $M_{Z}/2$.
It is immediately apparent that the $R_{pA}$ values (blue squares) align closely with the $r^{\textrm{A}}_{i}$ for light quarks across most of the kinematic region. However, at large $X_{B}$ values ($>0.1$), the $R_{pA}$ shows an excess, primarily due to an isospin effect. Specifically, the valence quark ratio $d/u$ in the nucleus is higher than that in the proton, and the stronger weak coupling of the $d$ quark leads to enhanced $Z$-boson production in nuclear collisions Halzen:2013bqa . This isospin effect can be quantified by setting all $r^{\textrm{A}}_{i}$ factors to unity in the calculation of $R_{pA}$, expressed schematically as
@@
Figure 4: Reorganized differential cross sections as functions of $X_{B}$, calculated at LO. Left panel: Results for $Z+$jet production, with contributions from nuclear quarks and gluons depicted as blue and red areas, respectively (inserted panel displays corresponding contribution fractions). Middle panel: Results for $Z$-boson production. Right panel: Rescaled results for $Z+$jet production, where nuclear quark contribution aligns with $Z$-boson cross section shown in middle panel. Calculations are performed at $\mu_{F}\!=\!\mu_{R}\!=\!\mu_{0}\!=\!M_{Z}/2$, using $pp$ collisions with backward-going proton treated as a nucleus.
@@ III A multi-process imaging of $r^{\textrm{A}}_{i}$
In this section, we extend the method of reorganizing the cross section to the study of $Z+$jet and $Z\!+\!c\!-\!$jet production. By leveraging the kinematic scan of nuclear PDFs in these processes, it becomes possible to separately enhance the signals of $r^{\textrm{A}}_{i}$ for gluon and charm quark distributions through a multi-process joint analysis.
For $Z+$jet production, the reorganized differential cross section can be defined in terms of $X_{B}$, $M_{V\!J}$, and $p_{T,avg}$ as outlined in Eq. (3). At LO, the initial state involves both quarks and gluons, meaning the $R_{pA}$ for such cross sections incorporates additional contributions from nuclear gluons compared to $Z$-boson production. In the left panel of Fig. 4, we present an example of the LO results for the reorganized cross section of $Z+$jet production in $p$A collisions. The contributions from nuclear quarks and gluons are depicted as blue and red areas, respectively. The inserted percentage diagram clearly illustrates that the contributions from nuclear gluons increase as $X_{B}$ decreases, consistent with expectations.
To isolate the nuclear gluon signal from the cross section, we assume that, under specific kinematic constraints, the nuclear modifications in quark-initiated processes approximate the $R_{pA}$ for $Z$-boson production at the LO level. This assumption will be validated later. Given that the differential cross section for $Z$-boson production (shown in the middle panel of Fig. 4) differs from the quark-initiated contribution to the $Z+$jet cross section, we introduce a ratio between the two as follows
@@
By multiplying the whole $Z+$jet cross section by the coefficient $\kappa_{1}(X_{B})$, we obtain a rescaled $Z+$jet cross section, as illustrated in the right panel of Fig. 4. At LO, the nuclear-quark component of the rescaled cross section aligns with the $Z$-boson production cross section. Further, by defining a difference term as
@@
$R_{\textrm{NLO}}(X_{B})\!=\!\frac{d\sigma^{Z}_{\textrm{NLO}}(X_{B})/d\sigma^{Z}_{\textrm{LO}}(X_{B})}{\,\,\,d\sigma^{Z\!+\!\textrm{jet}}_{\textrm{NLO}}(X_{B})/d\sigma^{Z\!+\!\textrm{jet}}_{\textrm{LO}}(X_{B})\,\,\,}.$ (8)
This factor accounts for NLO effects to some extent.
@@
Figure 5: Nuclear modification factors $R_{pA}$ as functions of $X_{B}$, calculated at NLO for reorganized cross section of $Z+$jet production (blue squares) and for combined observable $C_{g}(X_{B})$ (red discs). Predictions using EPPS21, nCTEQ15, and TUJU19 nuclear PDFs are displayed in left, middle, and right panels, respectively. Underlying $r^{\textrm{A}}_{i}(x,Q^{2})$ factors for light quarks and gluons from corresponding nuclear PDFs at $Q\!=\!M_{Z}/2$ are included in each panel for comparison. Bottom panels illustrate differences between each result and $r^{\textrm{A}}_{i}$ for gluon. NLO calculations are conducted at scale $\mu_{0}\!=\!M_{Z}/2$ for $Z$-boson and at $\mu_{0}\!=\!(M_{V\!J}\!+\!p_{T,avg})/4$ for $Z+$jet. Uncertainties in the results, evaluated by varying scale between $\mu_{0}/2$ and $2\mu_{0}$, are represented by shaded bands.
@@
In the three panels of Fig. 5, the NLO results for the $R_{pA}$ of individual $Z+$jet production, as defined in Eq. (4), are plotted as a function of $X_{B}$ (blue squares) using three sets of nuclear PDFs. By comparing these $R_{pA}$ values with the $r^{\textrm{A}}_{i}(x_{\textrm{Pb}})$ curves for light-quark (solid) and gluon (dotted-dashed) distributions in each panel, it is evident that the $R_{pA}$ squares generally lie between the quark and gluon curves. This reflects the fact that both nuclear parton species contribute significantly. Additionally, we present the corresponding NLO results for the $R_{pA}$ of $C_{g}(X_{B})$, defined in Eq. (7), as red discs in each panel of Fig. 5. These results demonstrate that the $R_{pA}$ of $C_{g}$ closely follows the $r^{\textrm{A}}_{i}$ curves for the gluon distribution. For further comparison, the differences between each result and the corresponding $r^{\textrm{A}}_{i}$ for gluon are also plotted in the bottom panels of Fig. 5.
@@
The above results are calculated for $Z+$jet production using the kinematic cuts $15066<M_{V\!J}<$ 150 GeV and $6040<p_{T,avg}<$ 60 GeV. These cuts ensure that the effective probing scales in this process are comparable to those in $Z$-boson production within the range $11666<M_{V}<$ 116 GeV. As a test, we compute the average value of the quantity $(M_{V\!J}+p_{T,avg})/2$ for $Z+$jet production and find it to be close to the $Z$-boson mass $M_{Z}$. Additionally, in an LO test we observe that the $R_{pA}$ of $C_{g}$ aligns well with the $r^{\textrm{A}}_{i}$ for the gluon distribution, indicating that the nuclear-quark-initiated processes in $Z+$jet production experience modifications similar to those in $Z$-boson production.
In the calculations, the renormalization and factorization scales are set to $\mu_{0}=(M_{V\!J}+p_{T,avg})/4$ for $Z+$jet production, while for $Z$-boson production, they are set to $\mu_{0}=M_{Z}/2$, consistent with the previous section. For $Z+$jet production, the variations in $R_{pA}$ with $\mu_{0}$ scaled by a factor of 2, represented as shaded bands around the squares in Fig. 5, are as small as those in $Z$-boson production. To illustrate the scale dependence for the $R_{pA}$ of the combined observable $C_{g}$, we assume that the choices of $\mu_{0}$ in the two processes are independent. Given that the ratio between $\mu^{2}_{0}$ in the two processes can vary by a factor of 16, the scale dependence of the combined observable generally appears stronger than that of a single observable, as shown in Fig. 5. The $R_{pA}$ of $C_{g}$ effectively images the $r^{\textrm{A}}_{i}$ for the gluon distribution up to a moderate value of $x_{\textrm{Pb}}$ ($\sim$0.1). For larger values of $x_{\textrm{Pb}}$, the uncertainties become significant, related to the rapidly diminishing contribution from gluons.

### v2

# Multiprocess imaging of nuclear modifications on parton distributions in proton-nucleus collisions

@@ Abstract
Nuclear modifications to collinear parton distribution functions are conventionally quantified by the ratios $r^{\textrm{A}}_{i}(x,Q^{2})=f^{\textrm{A,proton}}_{i}(x,Q^{2})/f^{\textrm{proton}}_{i}(x,Q^{2})$. For a given nucleus $A$, these ratios generally depend on the parton momentum fraction $x$, the probing scale $Q^{2}$, and the parton species $i$. Determining these dependencies relies on a global analysis of diverse experimental data. However, in realistic observables, these dependencies are intricately intertwined, making their extraction challenging. In this paper, we propose a novel approach to effectively image the nuclear modification factors $r^{\textrm{A}}_{i}(x,Q^{2})$ at the observable level in proton-nucleus collisions at the Large Hadron Collider. Specifically, through a combined study of $Z$-boson production, ZZ+jet production, and $Z+c$-jet production, we separately enhance signals arising from light-quark, gluon, and heavy-flavor (charm) distributions in nuclei. This enables us to effectively image the $r^{\textrm{A}}_{i}(x,Q^{2})$ for specific parton species. The feasibility of this method is validated through perturbative calculations at next-to-leading order in the strong coupling constant, employing three sets of nuclear PDF parametrizations: EPPS21, nCTEQ15, and TUJU19. Future measurements of these observables are expected to provide better-motivated parametrization form of nuclear PDFs and yield new insights into the detailed partonic structures of nuclei.
@@ I Introduction
Figure 1: Typical measurements of differential cross sections for $Z$-boson production (left panel), Z+Z\!+\!jet production (middle panel), and Z+cZ\!+\!c\!-\!jet production (right panel) in $pp$ collisions at $\sqrt{s}=8$ TeV at the LHC. Measurements are binned according to the rapidity $y^{Z}$ and transverse momentum $p_{T}^{Z}$ for $Z$-boson production CMS:2015hyl , the rapidity $y^{\textrm{jet}}$ and transverse momentum $p_{T}^{\textrm{jet}}$ for Z+Z\!+\!jet production ATLAS:2019bsa , and the transverse momentum $p_{T}^{\textrm{jet}}$ for Z+cZ\!+\!c\!-\!jet production CMS:2017snu . Experimental data are represented by black discs, while NLO predictions computed with MCFM Campbell:1999ah ; Campbell:2011bn ; Campbell:2015qma ; Campbell:2021vlt are depicted by red lines.
The vast amount of data from the Large Hadron Collider (LHC) has revolutionized global analyses over the past decade Klasen:2023uqj ; Klasen:2024xqn , leading to more stringent constraints on nuclear PDFs, notably for gluon distributions Kusina:2017gkz ; Eskola:2019dui ; Eskola:2019bgf ; Kusina:2020dki ; Paakkinen:2022qxn , and advancing neural-network-based analyses that reduce reliance on specific parametrizations AbdulKhalek:2022fyi . However, traditional measurements in proton-proton ($pp$) and proton-nucleus (ppA) collisions exhibit a less direct mapping to PDFs, and the imaging of $r^{\textrm{A}}_{i}(x,Q^{2})$ achieved in DIS has not been replicated for most LHC processes ALICE:2013snh ; LHCb:2017ygo ; CMS:2018jpl ; ALICE:2023ama ; ATLAS:2024mvt . This makes systematic comparisons of nuclear modifications across different processes challenging. Furthermore, significant discrepancies among widely used nuclear PDF parametrizations in certain regions remain unresolved Klasen:2023uqj ; Klasen:2024xqn .
In a previous study Shen:2021eir , a method to reorganize the differential cross section for dijet production, initially proposed by Ellis and Soper Ellis:1994dg , was extended to ppA collisions. This approach enables a well-controlled kinematic scan over $x$ and $Q^{2}$ for nuclear PDFs, providing an overall image of the $r^{\textrm{A}}_{i}(x,Q^{2})$ factors Shen:2021eir . However, this image combines contributions from different parton species simultaneously. Given that the relative contributions of various parton species vary across processes, extending this method to other ppA collision processes is of significant interest. Specifically, once kinematic scans are established for these processes, it becomes possible to enhance signals for specific parton flavors by combining results from different processes.
In this paper, we extend the method from Ref. Shen:2021eir to $Z$-boson, Z+Z+jet, and Z+Z+charm-jet production in proton-lead (ppPb) collisions at the LHC, enabling kinematic scans of nuclear PDFs for each process. The partonic subprocesses at leading order (LO) are listed in Tab. 1.
@@
Table 1: Partonic subprocesses at LO for $Z$-boson production, Z+Z+jet production, and Z+Z+charm-jet production, respectively.
@@
The paper is organized as follows: In Sec. II, we extend the method from Ref. Shen:2021eir to reorganize the differential cross sections for $Z$-boson, Z+Z+jet, and Z+Z+charm-jet production in ppPb collisions. We then present an imaging of $r^{\textrm{A}}_{i}(x,Q^{2})$ for light quarks using $Z$-boson production. In Sec. III, we design two combinations of multi-process cross sections and show that their nuclear modifications can image $r^{\textrm{A}}_{i}(x,Q^{2})$ for gluon and heavy-flavor (charm) distributions, respectively. A summary and discussion are provided in Sec. IV.
@@ II Reorganizing differential cross section in proton-nucleus collisions
Traditional measurements in $pp$ and ppA collisions are typically conducted in the form of differential cross sections, often double-differential, binned with variables such as the transverse momentum, rapidity, and invariant mass of the final-state particles Albacete:2013ei ; Albacete:2017qng . For instance, illustrated in Fig. 1 are three representative measurements for $Z$-boson production CMS:2015hyl , Z+Z+jet production ATLAS:2019bsa , and Z+Z+charm-jet productionCMS:2017snu in $pp$ collisions at the LHC, respectively. In general, such observables may not be ideal for probing the PDFs defined in terms of the parton momentum fraction $x$ and the probing scale $Q^{2}$. For example, in the case of Z+Z+jet production, measuring the cross section binned with the jet transverse momentum $p_{T}^{\textrm{jet}}$ and the jet rapidity $y^{\textrm{jet}}$ (as shown in the middle panel of Fig. 1) only constrains the kinematics of the jet, while leaving the kinematics of the $Z$ boson unconstrained. Consequently, even at LO, neither the momentum fraction carried by the initial-state parton nor the physical scale of the process can be precisely deduced from these two measured variables alone.
A well-controlled kinematic scan of the PDFs can be achieved by reorganizing the measured variables to establish a correspondence with $x$ and $Q^{2}$ at LO Shen:2021eir . To constrain the momentum fraction carried by the initial-state nuclear parton in ppA collisions (with the nucleus moving backward), a variable $X_{B}$ can be introduced Shen:2021eir ; Ellis:1994dg , generally expressed as
@@
Here, the summation of the backward components of the light-cone momenta, $p_{i}^{-}=E_{Ti}e^{-y_{i}}$, is performed over all inclusively produced particles. Specifically, for $Z$-boson production with leptonic decay ($Z\rightarrow l^{+}l^{-}$), the summation includes the di-lepton, while for Z+Z+jet or Z+cZ\!+\!c-jet production, it encompasses both the particles in the jet and the di-lepton from $Z$ decay. With this definition, $X_{B}$ precisely equals the momentum fraction $x_{\textrm{Pb}}$ of the parton from the lead nucleus in ppPb collisions at LO accuracy. However, this equivalence may not hold at NLO and beyond.
@@
For Z+(c)Z\!+\!(c)jet production, which involves more complex dynamic channels, one can employ both the invariant mass of the $Z$-jet system, $M_{V\!J}$, and the averaged transverse momentum, $p_{T,avg}$, to define a triple differential cross section:
@@
As an initial test, we investigate the $R_{pA}$ factor for the reorganized differential cross section of $Z$-boson production in ppPb collisions, as defined in Eq. (2). Since the partonic processes at LO are initiated by a $q\bar{q}$ pair, the $R_{pA}$ factors are expected to primarily reflect the nuclear modifications on the light-quark distributions. In the three panels of Fig. 3, the NLO results for $R_{pA}$ are plotted as a function of $X_{B}$ (blue squares) using three sets of nuclear PDFs. The variable $M_{V}$ is constrained around the $Z$-mass peak, specifically within the range $66<M_{V}<116$ GeV. To contextualize these modifications, the $r^{\textrm{A}}_{i}(x,Q^{2})$ curves for light-quark (solid) and gluon (dotted-dashed) distributions from the corresponding nuclear PDFs are also plotted in each panel of Fig. 3 as a function of $x_{\textrm{Pb}}$, with the probing scale $Q$ set at $M_{Z}/2$.
It is immediately apparent that the $R_{pA}$ values (blue squares) align closely with the $r^{\textrm{A}}_{i}$ for light quarks across most of the kinematic region. However, at large $X_{B}$ values (>0.1>0.1), the $R_{pA}$ shows an excess, primarily due to an isospin effect. Specifically, the valence quark ratio $d/u$ in the nucleus is higher than that in the proton, and the stronger weak coupling of the $d$ quark leads to enhanced $Z$-boson production in nuclear collisions Halzen:2013bqa . This isospin effect can be quantified by setting all $r^{\textrm{A}}_{i}$ factors to unity in the calculation of $R_{pA}$, expressed schematically as
@@
Figure 4: Reorganized differential cross sections as functions of $X_{B}$, calculated at LO. Left panel: Results for Z+Z+jet production, with contributions from nuclear quarks and gluons depicted as blue and red areas, respectively (inserted panel displays corresponding contribution fractions). Middle panel: Results for $Z$-boson production. Right panel: Rescaled results for Z+Z+jet production, where nuclear quark contribution aligns with $Z$-boson cross section shown in middle panel. Calculations are performed at $\mu_{F}\!=\!\mu_{R}\!=\!\mu_{0}\!=\!M_{Z}/2$, using pppp collisions with backward-going proton treated as a nucleus.
@@ III A multi-process imaging of $r^{\textrm{A}}_{i}$
In this section, we extend the method of reorganizing the cross section to the study of Z+Z+jet and Z+cZ\!+\!c\!-\!jet production. By leveraging the kinematic scan of nuclear PDFs in these processes, it becomes possible to separately enhance the signals of $r^{\textrm{A}}_{i}$ for gluon and charm quark distributions through a multi-process joint analysis.
For Z+Z+jet production, the reorganized differential cross section can be defined in terms of $X_{B}$, $M_{V\!J}$, and $p_{T,avg}$ as outlined in Eq. (3). At LO, the initial state involves both quarks and gluons, meaning the $R_{pA}$ for such cross sections incorporates additional contributions from nuclear gluons compared to $Z$-boson production. In the left panel of Fig. 4, we present an example of the LO results for the reorganized cross section of Z+Z+jet production in ppA collisions. The contributions from nuclear quarks and gluons are depicted as blue and red areas, respectively. The inserted percentage diagram clearly illustrates that the contributions from nuclear gluons increase as $X_{B}$ decreases, consistent with expectations.
To isolate the nuclear gluon signal from the cross section, we assume that, under specific kinematic constraints, the nuclear modifications in quark-initiated processes approximate the $R_{pA}$ for $Z$-boson production at the LO level. This assumption will be validated later. Given that the differential cross section for $Z$-boson production (shown in the middle panel of Fig. 4) differs from the quark-initiated contribution to the Z+Z+jet cross section, we introduce a ratio between the two as follows
@@
By multiplying the whole Z+Z+jet cross section by the coefficient $\kappa_{1}(X_{B})$, we obtain a rescaled Z+Z+jet cross section, as illustrated in the right panel of Fig. 4. At LO, the nuclear-quark component of the rescaled cross section aligns with the $Z$-boson production cross section. Further, by defining a difference term as
@@
$R_{\textrm{NLO}}(X_{B})\!=\!\frac{d\sigma^{Z}_{\textrm{NLO}}(X_{B})/d\sigma^{Z}_{\textrm{LO}}(X_{B})}{\,\,\,d\sigma^{Z\!+\!\textrm{jet}}_{\textrm{NLO}}(X_{B})/d\sigma^{Z\!+\!\textrm{jet}}_{\textrm{LO}}(X_{B})\,\,\,}$, (8)
which takes into account the differences between the NLO K-factor effects in the two combined processes to some extent.
@@
Figure 5: Nuclear modification factors $R_{pA}$ as functions of $X_{B}$, calculated at NLO for reorganized cross section of Z+Z+jet production (blue squares) and for combined observable $C_{g}(X_{B})$ (red discs). Predictions using EPPS21, nCTEQ15, and TUJU19 nuclear PDFs are displayed in left, middle, and right panels, respectively. Underlying $r^{\textrm{A}}_{i}(x,Q^{2})$ factors for light quarks and gluons from corresponding nuclear PDFs at $Q\!=\!M_{Z}/2$ are included in each panel for comparison. Bottom panels illustrate differences between each result and $r^{\textrm{A}}_{i}$ for gluon. NLO calculations are conducted at scale $\mu_{0}\!=\!M_{Z}/2$ for $Z$-boson and at $\mu_{0}\!=\!(M_{V\!J}\!+\!p_{T,avg})/4$ for Z+Z+jet. Uncertainties in the results, evaluated by varying scale between $\mu_{0}/2$ and $2\mu_{0}$, are represented by shaded bands.
@@
In the three panels of Fig. 5, the NLO results for the $R_{pA}$ of individual Z+Z+jet production, as defined in Eq. (4), are plotted as a function of $X_{B}$ (blue squares) using three sets of nuclear PDFs. By comparing these $R_{pA}$ values with the $r^{\textrm{A}}_{i}(x_{\textrm{Pb}})$ curves for light-quark (solid) and gluon (dotted-dashed) distributions in each panel, it is evident that the $R_{pA}$ squares generally lie between the quark and gluon curves. This reflects the fact that both nuclear parton species contribute significantly. Additionally, we present the corresponding NLO results for the $R_{pA}$ of $C_{g}(X_{B})$, defined in Eq. (7), as red discs in each panel of Fig. 5. These results demonstrate that the $R_{pA}$ of $C_{g}$ closely follows the $r^{\textrm{A}}_{i}$ curves for the gluon distribution. For further comparison, the differences between each result and the corresponding $r^{\textrm{A}}_{i}$ for gluon are also plotted in the bottom panels of Fig. 5.
@@
The above results are calculated for Z+Z+jet production using the kinematic cuts $66<M_{V\!J}<150$ GeV and $40<p_{T,avg}<60$ GeV. These cuts ensure that the effective probing scales in this process are comparable to those in $Z$-boson production within the range $66<M_{V}<116$ GeV. As a test, we compute the average value of the quantity $(M_{V\!J}+p_{T,avg})/2$ for Z+Z+jet production and find it to be close to the $Z$-boson mass $M_{Z}$. Additionally, in an LO test we observe that the $R_{pA}$ of $C_{g}$ aligns well with the $r^{\textrm{A}}_{i}$ for the gluon distribution, indicating that the nuclear-quark-initiated processes in Z+Z+jet production experience modifications similar to those in $Z$-boson production.